The sides of a quadrilateral are all positive integers and three of them are 30,80 and 90 units.How many possible values are there for the fourth side?

Grade:10

1 Answers

Arun
5 years ago
property of quadrilateralis all sides of quadrilaterial is less than sum of all other sides
sides are of length
30, 80, 90
so maximum length can be, 30+80+90-1= 199 ( as it's less than not equal to sum of all other sides)
and minimum can be, let it be x
90=80+ 30 +x
=>x=20
x=21 ( as it's less than not equal to sum of all other sides)
so, from 21 to 199 interger can be
199- 20 = 179 answer